OpenStudy (anonymous):

Solve the system of inequalities by using the substitution method. Check each solution. Show all your work. y ≤ x – 2 x + 3y ≥ 6 Solve the system of equations by using the substitution method. Check each solution. Show all your work. 3x + 2y = 6 x + 4y = -8

OpenStudy (anonymous):

ohhh not sure on the first one but for the second one i rearranged them so the equations looked like this: 4y=x+8 2y=3x-6 can you see how i got that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i would then try to isolate one of the Ys. 4y=x+8 would turn into y=x+2. then plug it in. 2(x+2)=3x-6 do distruibutive property 2x+4=3x-6 -2x -2x 4=x-6 +6 +6 10=x

OpenStudy (anonymous):

OK!

OpenStudy (anonymous):

Then plug in the answer to any equation you want. you got it from there?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

Wait, what do you mean by pluf in the answer int o any equation?

OpenStudy (anonymous):

I mena plug in the answer to any equaiton

OpenStudy (anonymous):

So like 3(10) + 2y = 6 Y wasn't found?